A company issues $25750000, 7.8%, 20-year bonds to yield 8% on January 1, 2020. Interest is paid on June 30 and December 31. The proceeds from the bonds are $25240335. Using effective-interest amortization, what will the carrying value of the bonds be on the December 31, 2020 balance sheet? (Round answer to 0 decimal place, e.g. 52.)


a. $25750000

b. $25245696

c. $25266879

d. $25251276
Answer #1

d. $25251276

Semi annual ending Beginning Carrying Value Interest Expense Coupon Interest Reduction of principal Ending Carrying Value
a b=a*4% c d=b-c e=a-d
June 30, 2020 25240335 1009613 1004250 5363.4 25245698
December 31, 2020 25245698 1009828 1004250 5577.936 25251276
Working:
Semi annual coupon = 25750000 * 0.039
= 1004250